OpenStudy (anonymous):

the curve y^2=x(x-n)^2, when n is a psitive integer,encloses a loop. find exact area enclosed by this loop for n=1,2,3 and 4

OpenStudy (anonymous):

try: \[\int\limits_{0}^{n} -2\sqrt{x}(x-n) dx\]

OpenStudy (anonymous):

Above times two.

OpenStudy (anonymous):

why's that?

OpenStudy (anonymous):

...shoot it was y^2=x^2(x-n)

OpenStudy (anonymous):

@Algebraic! It only gives area with x axis. The area was both upwards and downwards.

OpenStudy (anonymous):

sqyaure root of x^2 is x

OpenStudy (anonymous):

y= +/- sqrt(x)(x-n)

OpenStudy (anonymous):

nice sid

OpenStudy (anonymous):

@Algebraic! Even if its +- You are integrating it just once right?

OpenStudy (anonymous):

I thought the y^2 cancxeled out the x^2?

OpenStudy (anonymous):

you integrate the top portion, since it's positive, then you double it.

OpenStudy (anonymous):

so...Its x square root of x-n)?

OpenStudy (anonymous):

or, alternately, integrate the bottom portion times -2

OpenStudy (anonymous):

Precisely.

OpenStudy (anonymous):

l2read

OpenStudy (anonymous):

ok... the formula for area as a function of n...?\[A=2\int\limits_{0}^{?}\]

OpenStudy (anonymous):

\[A=\int\limits_{0}^{n}-2x \sqrt{(x-n)}\]

OpenStudy (anonymous):

??

OpenStudy (phi):

note: you said ..shoot it was y^2=x^2(x-n) But this does not form a loop. see http://www.wolframalpha.com/input/?i=y%5E2%3Dx%5E2%28x-n%29%2C+where+n%3D2 On the other hand, the original equation does form a loop http://www.wolframalpha.com/input/?i=y%5E2%3Dx%28x-n%29%5E2%2C+where+n%3D2

OpenStudy (anonymous):

you are right